sland Novelties, Inc., of Palau makes two products—Hawaiian Fantasy and Tahitian Joy. Each product's selling price, variable expense per unit and annual sales volume are as follows:

Hawaiian Fantasy Tahitian Joy
Selling price per unit $ 20 $ 110
Variable expense per unit $ 9 $ 33
Number of units sold annually 22,000 6,000

Fixed expenses total $664,000 per year.

Required:

1. Assuming the sales mix given above, do the following:

a. Prepare a contribution format income statement showing both dollar and percent columns for each product and for the company as a whole.

b. Compute the company's break-even point in dollar sales. Also, compute its margin of safety in dollars and its margin of safety percentage.

2. The company has developed a new product called Samoan Delight that sells for $30 each and that has variable expenses of $24 per unit. If the company can sell 22,000 units of Samoan Delight without incurring any additional fixed expenses:

a. Prepare a revised contribution format income statement that includes Samoan Delight. Assume that sales of the other two products does not change.

b. Compute the company’s revised break-even point in dollar sales. Also, compute its revised margin of safety in dollars and margin of safety percentage.
Answer #1
1) Hawaiian Fantasy Tahitian Joy Total
a) % % %
Sales (22000*20/6000*110) 440000 100.00% 660000 100.00% 1100000 100.00%
Variable cost (22000*9+6000*33) 198000 45.00% 198000 30.00% 396000 36.00%
Contribution margin 242000 55.00% 462000 70.00% 704000 64.00%
Fixed expenses 664000 60.36%
Net operating income 40000 3.64%
b) BEP in dollars = Fixed cost/CM ratio = 664000/64% = 1037500
Margin of safety in dollars = Current sales-BEP sales = 1100000-1037500 = 62500
Margin of safety in %= (Current sales-BEP sales)/Current sales = (1100000-1037500)/1100000 = 5.68%
2)
a) Hawaiian Fantasy Tahitian Joy Samoan Delight Total
% % % %
Sales (22000*20/6000*110/22000*30) 440000 100.00% 660000 100.00% 660000 100.00% 1760000 100.00%
Variable cost (22000*9+6000*33) 198000 45.00% 198000 30.00% 528000 80.00% 924000 52.50%
Contribution margin 242000 55.00% 462000 70.00% 132000 20.00% 836000 47.50%
Fixed expenses 664000 37.73%
Net operating income 172000 9.77%
BEP in dollars = Fixed cost/CM ratio = 664000/47.50% = 1397895
Margin of safety in dollars = Current sales-BEP sales = 1760000-1397895 = 362105
Margin of safety in %= (Current sales-BEP sales)/Current sales = (1760000-1397895)/1760000 = 20.57%